if you draw 2 clubs on 2 consecutive draws from a standard deck of cards you win $23 otherwise you pay me $3 cards drawn without replacement enter the expected value of proposition
Answer by edjones(8007)

13/52 * 12/51 = 1/17
16*-3=-48
23*1 =23
----------add
=====>-$25. expected value of proposition.
.
Ed